# Not All Tokens Are Meant to Be Forgotten

### v1

@@ 4.1 Datasets and Metrics
(1) MUSE is a benchmark for unlearning the copyrighted content with two unlearning tasks: forgetting the Harry Potter books (termed ‘Books’) and news articles (termed ‘News’), respectively. To evaluate the effectiveness of unlearning and the preservation of utility for MUSE, we use three metrics: Verbatim Memorization (VerbMem), Knowledge Memorization (KnowMem), and Privacy Leakage (PrivLeak). VerbMem and KnowMem are measured using ROUGE-L F1 rouge2004package , where lower scores indicate reduced verbatim and factual memorization, respectively. PrivLeak quantifies privacy risks using the Min-KProb metric shi2023detecting in a membership inference attack. A value close to zero indicates minimal privacy leakage, while large positive/negative values suggest over/under-forgetting. We conduct our experiments on MUSE using ICLM-7B shi2023context and LLaMA-2 7B touvron2023llama .
Benchmark Used LLM Forget quality Model Utility
MUSE ICLM-7B KnowMem on $\mathcal{D}_{f}$downarrow KnowMem on $\mathcal{D}_{r}\quad\uparrow$
VerbMem on $\mathcal{D}_{f}$downarrow
LLaMa-2 7B PrivLeak $(\rightarrow 0)$
TOFU LLaMa-2 7B Truth Ratio on $\mathcal{D}_{f}$uparrow $\text{Mean}\Bigg{(}\begin{cases}\text{Probability, Rouge-L, Truth Ratio}\\ \mathcal{D}_{r},\mathcal{D}_{\text{real\_authors}},\mathcal{D}_{\text{word\_facts}}\end{cases}\!\!\!\!\!\!\!\Bigg{)}$uparrow
LLaMa-3.2 3B
@@
(2) TOFU is a synthetic Q&A dataset of 200 author biographies with three unlearning tasks: forget 1%, 5%, and 10% of the author profiles. We evaluate the unlearning performance for TOFU using two key metrics: Forget Quality and Model Utility as defined in maini2024tofu . Forget quality is quantified using the $p$-value from a Kolmogorov-Smirnov (KS) test, where a higher $p$-value indicates greater similarity between the output distributions of the unlearned and the retained model. The retained model here refers to retraining an LLM from scratch on the retain dataset while excluding the forget set and is regarded as the gold standard for unlearning maini2024tofu ; zhang2024negative . Model utility measures the model’s performance on the retain set and its ability to retain real-world knowledge. This is assessed using various metrics, including ROUGE-L rouge2004package and Truth Ratio maini2024tofu . Our experiments on TOFU utilize LLaMA-2 7B and LLaMA-3.2 3B dubey2024llama .
@@
The LLM models and the evaluation metrics used across various unlearning benchmarks are summarized in Table 1.
4.2 Unlearning Baselines
@@
We compare our method with baselines, i.e., GA, NPO, and SimNPO, on both MUSE and TOFU. For other baselines, such as Task Vector for MUSE and Kahneman-Tversky Optimization (KTO) for TOFU, we strictly follow their original implementations outlined in their respective benchmarks. We also evaluate the impact of incorporating Gradient Descent on the retain (GDR) loss with the baselines, i.e., GAGDR, NPOGDR, SimNPOGDR, and TPOGDR, on MUSE. Specifically, the GDR loss liu2022continual ; yao2023large ; zhang2024negative ; shi2024muse is a standard gradient descent objective applied to the cross-entropy loss on the retain set $\mathcal{D}_{r}$. This approach enables the model to be explicitly trained to maintain performance on the retain set $\mathcal{D}_{r}$. More details of all baseline methods are provided in Appendix D.
@@
4.3 Unwanted Information Identifier
We employ two different unwanted information identifiers for TOFU dataset: a generative LM using ChatGPT-4o (via the web interface) and a discriminative LM using DistilBERT (Section 3.1.1). In Appendix C, we further examine the effectiveness of unlearning methods using both identifiers, showing that the generative LM approach enables a better balance between forget quality and model utility compared to the discriminative LM approach. Each sample in the forget set of the MUSE Books dataset contains approximately 175k words (more than 200k tokens), whereas current GPT models, including ChatGPT-4o, can only handle a maximum token window size of 128k tokens (roughly 100k words). Therefore, the GPT models cannot process all the information from individual samples. It is challenging to achieve stable and consistent UW identification with GPT models for the MUSE dataset. We thus only adopt the discriminative LM approach as the unwanted information identifier on this dataset.
Computational Efficiency: Distinguishing between UW from GW using either generative or discriminative approaches remains computationally efficient and time-effective. For smaller data sets like TOFU, using a generative LM like ChatGPT-4o to process it takes several minutes. For larger datasets like MUSE, using a discriminative model like DistilBERT on our H100 server completes the task in just a few hours. These customized approaches highlight our method’s flexibility, computational efficiency, and scalability across diverse dataset sizes.
5 Results and Discussion
5.1 Performance on MUSE
GDR significantly improves the utility preservation. As shown in the Table 2, nearly all unlearning methods (except Task Vector) suffer from severe utility degradation on the MUSE benchmark when GDR is not used. This is largely due to the large forget set size of the MUSE Benchmark. In contrast, incorporating GDR consistently improves utility across all methods. However, Task Vector fails completely in unlearning on both MUSE Books and News, as its forget quality remains nearly identical to that of the original model.
TPOGDR consistently achieve optimal PribLeak values. PrivLeak in the results serves as the primary metric to measure the performance gap with the retained model defined in Section 1. Notably, our TPOGDR method consistently achieves PrivLeak values closest to 0 on both News (-6.12) and Books (-21.75), while also attaining comparable KnowMem and VerbMem scores on the forget set relative to other GDR-based baselines. Additionally, TPO achieves comparable model utility to the strongest baseline, SimNPOGDR, while surpassing it in forget quality. This highlights TPO’s effectiveness in unlearning and achieve superior unlearning performance compared to other baselines
5.2 Performance on TOFU
@@
TPO-GPT achieves the best forget quality on a larger forget set size. Figure 6 illustrates that all baseline methods experience a significant decline in forget quality as the forget set size increases. Notably, at a forget set size of 10%, GA-based and KTO-based methods fail completely in unlearning for both LLaMA2 7B-Chat and LLaMA3.2 3B models, evidenced by their near-zero forget quality. Further, while NPO-based and SimNPO-based methods achieve higher forget quality, their performance also noticeably declines when the forget set size reaches 10%. In contrast, the developed TPO-GPT consistently demonstrates comparable forget quality on smaller forget set sizes (e.g., 1% and 5%) and achieves optimal forget quality on larger forget set sizes (e.g., 10 %) for both models. Notably, on LLaMA2 7B-Chat, TPO-GPT consistently achieves forget quality exceeding 90% across different forget set sizes, as evidenced by the first row of Figure 6.
@@
6 Related Work
@@
LLM Unlearning. Motivated by data privacy regulations like the General Data Protection Regulation (GDPR) that gave individual users the “right to be forgotte” rosen2011right , machine unlearning was initially developed to remove the effect of specific training examples without retraining the model on entire data cao2015towards ; bourtoule2021machine . The effectiveness of machine unlearning has been shown in different domains including image classification sekhari2021remember ; fan2025challenging and federated learning wang2022federated ; jin2023forgettable . However, these unlearning methods often become in-scalable for LLMs due to massive parameter sizes in LLMs.
@@
Recent advancements have sought to extend machine unlearning techniques to LLMs. The majority of works focus on fine-tuning the model by gradient ascent on the loss computed on the forget set, alongside gradient descent or KL divergence on the loss computed on the retain set li2024wmdp ; yao2023large ; chen2023unlearn ; jang2022knowledge ; wang2023kga ; tian2024forget ; liu2024towards ; ji2024reversing ; wang2024llm ; zhao2024makes . Despite these developments, existing methods yao2023large ; liu2022continual frequently struggle to balance between forgetting quality and utility preservation, frequently resulting in catastrophic collapse, as observed on benchmark datasets like TOFU for fictitious unlearning maini2024tofu . To address this limitation, NPO zhang2024negative , inspired by direct preference optimization (DPO) rafailov2024direct , introduces a lower-bounded unlearning objective to mitigate catastrophic collapse. Additionally, Simple Negative Preference Optimization (SimNPO) fan2024simplicity enhances NPO by proposing a reference-free variant, drawing inspiration from Simple Preference Optimization (SimPO) meng2024simpo . Additionally, SOUL jia2024soul demonstrates improved unlearning efficacy by leveraging second-order optimization. However, the performance of these methods deteriorates as the size of the forget set increases liu2024large , underscoring the need for more robust and scalable solutions to achieve effective unlearning while preserving model utility.
@@
Targeted Unlearning. Recent work, such as RESTOR rezaei2024restor , demonstrates that isolating and precisely targeting the information within the unlearning scope significantly improves the unlearning performance of existing methods (e.g., GA yao2023large ), highlighting the crucial role of targeted unlearning. However, a major challenge remains: knowledge dependencies liu2024large make it difficult to cleanly separate the information that should be forgotten from what should be retained. Recent methods tackle this challenge through varied strategies: MemFlex tian2024forget leverages gradient information to focus on sensitive parameters accurately. ECO liu2024large proposes an efficient unlearning framework that localizes unlearning to sentences that contain content within the unlearning target by using a sentence-level identifier. However thaker2024position shows that ECO’s classifier tends to overfit specific keywords rather than aligning with the unlearning requester’s intent, leading to insufficient unlearning when those keywords are absent or rephrased in the sentence. Additionally, SEUL wang2024selective achieves fine-grained information identification by incorporating a sensitive span annotation framework that uses a LLM (e.g., ChatGPT) to annotate specific spans containing sensitive information. This approach improves unlearning effectiveness by targeting specific continuous sequence spans. However, it focuses solely on PII unlearning, overlooking broader generalizability to various unlearning tasks. (e.g., copyrighted content unlearning). Despite these advancements, challenges such as over and under-forgetting remain, highlighting the need for more precise and robust solutions to disentangle information dependencies between forget and retention sets thaker2024position ; liu2024rethinking ; zhao2024makes .
7 Conclusion
@@
In this work, we introduced TIF, a framework designed to improve unlearning in LLMs by distinguishing between UW and GW. At its core, TIF incorporates the TPO objective, which unlearns UW and preserves GW. Experiments on TOFU and MUSE benchmarks demonstrate that TIF not only enhances unlearning effectiveness for existing unlearning methods but also preserves significantly more model utility. Here we focus on sequence unlearning via suppressing token generation, and demonstrate its applications to copyright or privacy protection. Yet, there is another line of work in knowledge unlearning (e.g., WMDP li2024wmdp ), which focuses on unlearning an entire distribution of hazardous knowledge embedded in the latent layer representations. These works have demonstrated effectiveness in enhancing LLM security, such as biosecurity, cybersecurity, and chemical security.
@@ Appendix A Appendix: Additional Experiment Details
Computational configurations. All experiments are conducted on 2 NVIDIA H100 GPU cards in a single node.
@@
For unlearning with NPO, we use the default setting, fixing the parameter $\beta$ at 0.1. For TPO, the parameter $\beta$ is tuned by searching within the range [0.1, 0.5] to obtain the best-performing model. We report the value of $\beta$ that yielded the best unlearning performance when TPO achieves the best forget quality across different tasks and models in Table 4.
@@
Appendix B Unlearning performance combining the GDR with the proposed TIF framework.
Table 5 summarizes the forget quality and model utility of our TPO compared to several baseline approaches (i.e., GA, KTO, NPO, SimNPO), evaluated on the TOFU Forget 05 task. The experiments are conducted under two conditions: with and without incorporating Gradient Descent on Retain (GDR) loss, and with and without our GPT-based unwanted information identifier (GPT). Here, "Vanilla" denotes methods evaluated without incorporating the GPT-based identifier.
@@
Appendix C Generative LM Approach vs Discriminative LM Approach.

### v2

@@ Introduction
(1) MUSE is a benchmark for unlearning the copyrighted content with two unlearning tasks: forgetting the Harry Potter books (termed ‘Books’) and news articles (termed ‘News’), respectively. To evaluate the effectiveness of unlearning and the preservation of utility for MUSE, we use three metrics: Verbatim Memorization (VerbMem), Knowledge Memorization (KnowMem), and Privacy Leakage (PrivLeak). VerbMem and KnowMem are measured using ROUGE-L F1 (rouge2004package), where lower scores indicate reduced verbatim and factual memorization, respectively. PrivLeak quantifies privacy risks using the Min-K% Prob metric (shi2023detecting) in a membership inference attack. A value close to zero indicates minimal privacy leakage, while large positive/negative values suggest over-/under-forgetting. We conduct our experiments on MUSE using ICLM-7B (shi2023context) and LLaMA-2 7B (touvron2023llama).
@@
(2) TOFU is a synthetic Q&A dataset of 200 author biographies with three unlearning tasks: forget 1%, 5%, and 10% of the author profiles. We evaluate unlearning performance using two key metrics: Forget Quality and Model Utility as defined in (maini2024tofu). Forget quality is quantified using the $p$-value from a Kolmogorov-Smirnov (KS) test, where a higher $p$-value indicates greater similarity between the output distributions of the unlearned and the retained model. The retained model denotes retraining an LLM from scratch on the retain dataset while excluding the forget set and is regarded as the gold standard for unlearning (maini2024tofu; zhang2024negative). Model utility measures the model’s performance on the retain set and its ability to retain real-world knowledge. This is assessed using various metrics, including ROUGE-L (rouge2004package) and Truth Ratio (maini2024tofu). Experiments on TOFU utilize LLaMA-2 7B and LLaMA-3.2 3B (dubey2024llama).
@@
The LLMs and the evaluation metrics across unlearning benchmarks are summarized in Table 5 (Appendix).
Unlearning Baselines
@@
We compare our method with baselines, i.e., GA, NPO, and SimNPO, on both MUSE and TOFU. For other baselines, such as Task Vector for MUSE and Kahneman-Tversky Optimization (KTO) for TOFU, we strictly follow their original implementations outlined in their respective benchmarks. We also evaluate the impact of incorporating Gradient Descent on the retain (GDR) loss with the baselines, i.e., GAGDR, NPOGDR, SimNPOGDR, and TPOGDR, on MUSE. Specifically, the GDR loss (liu2022continual; yao2023large; zhang2024negative; shi2024muse) is a standard gradient descent objective applied to the cross-entropy loss on the retain set $\mathcal{D}_{r}$. This approach enables the model to be explicitly trained to maintain performance on the retain set $\mathcal{D}_{r}$. More details of all baseline methods are provided in Appendix E.
Unwanted Information Identifier
We employ two different unwanted information identifiers for TOFU dataset: a generative LM using ChatGPT-4o (via the web interface) and a discriminative LM using DistilBERT (Section Discriminative Encoder-Only LM.). In Appendix D, we further examine the effectiveness of unlearning methods using both identifiers, showing that the generative LM approach enables a better balance between forget quality and model utility compared to the discriminative LM approach. For the MUSE dataset, which is non–QA in nature, we treat the beginning-of-sequence token <bos> as the input $x$, and define the remaining tokens as the target sequence $y$. Because each sample in the forget set of the MUSE Books dataset contains approximately 175k words (more than 200k tokens), whereas current GPT models, including ChatGPT-4o, can only handle a maximum token window size of 128k tokens (roughly 100k words). Therefore, the GPT models cannot process all the information from individual samples. It is challenging to achieve stable and consistent UW identification with GPT models for the MUSE dataset. We thus only adopt the discriminative LM approach as the unwanted information identifier on this dataset.
@@
Results and Discussion
Performance on TOFU
@@
TPO-GPT achieves the best forget quality on a larger forget set size. Figure 9 illustrates that all baseline methods experience a significant decline in forget quality as the forget set size increases. Notably, at a forget set size of 10%, GA-based and KTO-based methods fail completely in unlearning for both LLaMA2 7B-Chat and LLaMA3.2 3B models, evidenced by their near-zero forget quality. Further, while NPO-based and SimNPO-based methods achieve higher forget quality, their performance also noticeably declines when the forget set size reaches 10%.
In contrast, the developed TPO-GPT consistently demonstrates comparable forget quality on smaller forget set sizes (e.g., 1% and 5%) and achieves optimal forget quality on larger forget set sizes (e.g., 10 %) for both models. Notably, on LLaMA2 7B-Chat, TPO-GPT consistently achieves forget quality exceeding 90% across different forget set sizes, as evidenced by the first row of Figure 9.
@@
Performance on MUSE
GDR significantly improves the utility preservation. As shown in the Table 1, nearly all unlearning methods suffer from severe utility degradation on the MUSE benchmark when GDR is not used. This is largely attributed to the large size of the forget set in the MUSE Benchmark, which poses a challenge to preserving general model performance. Incorporating GDR consistently mitigates this issue and improves utility across all methods.
TPOGDR consistently achieves the best PrivLeak performance. PrivLeak serves as the primary metric to measure the performance gap from the retained model defined in Section Datasets and Metrics. Notably, our TPOGDR method consistently achieves PrivLeak values closest to 0 on both News (-6.12) and Books (-20.66), while maintaining comparable KnowMem and VerbMem scores on the forget set relative to other GDR-based baselines. Although Task Vector preserves high model utility on both tasks, it fails completely in unlearning, as its forget quality remains nearly identical to that of the original model. Overall, TPO offers the best trade-off between effective unlearning (lowest PrivLeak) and reasonable utility retention, showing superior performance over all baselines.
Related Work
@@
LLM Unlearning. Motivated by data privacy regulations like the General Data Protection Regulation (GDPR) that gave individual users the “right to be forgotten” (rosen2011right), machine unlearning was initially developed to remove the effect of specific training examples without retraining the model on the entire data (cao2015towards; bourtoule2021machine). Its effectiveness has been shown in different domains, including image classification (sekhari2021remember; fan2025challenging) and federated learning (wang2022federated; jin2023forgettable). However, these unlearning methods often become infeasible for LLMs due to the massive parameter sizes in LLMs.
@@
Recent efforts have adapted unlearning to LLMs by fine-tuning with gradient ascent on the forget set and gradient descent or KL divergence on the retain set (li2024wmdp; yao2023large; chen2023unlearn; jang2022knowledge; wang2023kga; tian2024forget; liu2024towards; ji2024reversing; wang2024llm; zhao2024makes; jia2024soul). Yet, existing methods (yao2023large; liu2022continual) often struggle to balance forgetting and utility preservation, frequently resulting in catastrophic collapse, as observed on benchmarks like TOFU (maini2024tofu). To address this limitation, NPO (zhang2024negative), inspired by direct preference optimization (DPO) (rafailov2024direct), introduces a lower-bounded unlearning objective to mitigate catastrophic collapse. Additionally, Simple Negative Preference Optimization (SimNPO) (fan2024simplicity) enhances NPO by proposing a reference-free variant, drawing inspiration from Simple Preference Optimization (SimPO) (meng2024simpo). However, the performance of these methods deteriorates as the size of the forget set increases (liu2024large), underscoring the need for more robust and scalable solutions to achieve effective unlearning while preserving model utility.
@@
Targeted Unlearning. Recent work, such as RESTOR (wan2025not; liu2024revisiting; rezaei2024restor), demonstrates that isolating and precisely targeting the information within the unlearning scope significantly improves the unlearning performance of existing methods (e.g., GA (yao2023large), WHP (eldan2023s)), highlighting the crucial role of targeted unlearning. However, a major challenge remains: knowledge dependencies (liu2024large) make it difficult to cleanly separate the information that should be forgotten from what should be retained. Recent methods tackle this challenge through varied strategies: MemFlex (tian2024forget) leverages gradient information to focus on sensitive parameters accurately. ECO (liu2024large) proposes an efficient unlearning framework that localizes unlearning to sentences that contain content within the unlearning target by using a sentence-level identifier. However, (thaker2024position) shows that ECO’s classifier tends to overfit specific keywords rather than align with the unlearning requester’s intent, leading to insufficient unlearning when those keywords are absent or rephrased in the sentence. Additionally, SEUL (wang2024selective) achieves fine-grained information identification by incorporating a sensitive span annotation framework that uses an LLM (e.g., ChatGPT) to annotate specific spans containing sensitive information. This approach improves unlearning effectiveness by targeting specific continuous sequence spans. However, it focuses solely on PII unlearning, overlooking broader generalizability to various unlearning tasks. (e.g., copyrighted content unlearning). Despite these advancements, challenges such as over- and under-forgetting remain, highlighting the need for more precise and robust solutions to disentangle information dependencies between forget and retention sets (thaker2024position; liu2024rethinking; zhao2024makes).
Conclusion
@@
In this work, we propose TIF, a framework that improves LLM unlearning by distinguishing between UW and GW. TIF employs the TPO objective to selectively unlearn UW while preserving GW. Experiments on TOFU and MUSE benchmarks show that TIF enhances unlearning effectiveness for existing unlearning methods and substantially preserves more model utility. Our study focuses on sequence unlearning by suppressing token generation, relevant to copyright and privacy protection. In contrast, knowledge unlearning (e.g., WMDP (li2024wmdp)) targets unlearning entire distributions of hazardous knowledge from latent representations, advancing model security in domains such as biosecurity, cybersecurity, and chemical safety.
Ethical Statement
This work develops methods for targeted unlearning in large language models to enhance privacy and compliance with data protection laws. All experiments use public datasets (TOFU and MUSE) without personal or sensitive data. We highlight that unlearning techniques should be applied responsibly to prevent misuse or selective information removal.
Acknowledgments
This work was supported in part by the National Science Foundation under Awards 2504264, 2211897, 2211896, and 2401775.
@@ Appendix A Appendix: Additional Experiment Details
A.1 Computational Configurations
All experiments are conducted on 2 NVIDIA H100 GPU cards in a single node.
@@ A.1 Computational Configurations
For unlearning with NPO, we use the default setting, fixing the parameter $\beta$ at 0.1. For TPO, the parameter $\beta$ is tuned by searching within the range [0.1, 0.5] to obtain the best-performing model. We report the value of $\beta$ that yielded the best unlearning performance when TPO achieves the best forget quality across different tasks and models in
A.2 Computational Efficiency of the Unwanted Information Identification
Distinguishing between UW from GW using either generative or discriminative approaches remains computationally efficient and time-effective. For smaller data sets like TOFU, using a generative LM like ChatGPT-4o to process it takes several minutes. For larger datasets like MUSE, using a discriminative model like DistilBERT on our H100 server completes the task in just a few hours. These customized approaches highlight our method’s flexibility, computational efficiency, and scalability across diverse dataset sizes.
A.3 Evaluation Metrics
We summarize the LLM models and the evaluation metrics used across various unlearning benchmarks in Table 5.
A.4 Training Cost Analysis
We evaluate the computational overhead of our proposed TPO-GPT compared to the baseline NPO-GPT. TPO-GPT requires only about $1.7\times$ more training time and incurs roughly $6\%$ higher GPU memory consumption. These results indicate that TPO introduces moderate additional cost while providing substantial unlearning benefits.
Appendix B Additional study on individual components of TPO
Our TPO loss comprises two components: the LDL and the PL. To further disentangle their individual contributions, we conduct an ablation study on the Forget-05 subset of TOFU. The role of PL in maintaining model utility has been discussed in Section Targeted Preference Optimization (TPO), and its effectiveness is further demonstrated by the results in Table 4. Both LDL and NPO exhibit reduced utility degradation when combined with the PL loss. Moreover, LDL plays a central role in TPO by effectively balancing forget quality and model utility. Compared to NPO variants, LDL alone preserves substantially more utility, and when integrated with PL and our GPT-based unwanted information identification process, it achieves the best trade-off between FQ and MU. These findings underscore the complementary roles of LDL and PL in enabling targeted unlearning.
@@ Appendix B Additional study on individual components of TPO
Forget 05 NPO NPO-GPT NPO+PL-GPT LDL-only LDL-GPT TPO (LDL+PL)-GPT
FQ 0.96 0.96 0.96 0.55 0.92 0.96
MU 0.11 0.21 0.42 0.51 0.63 0.67
Table 4: Comparison of unlearning performance on the Forget-05 subset of TOFU between LDL (a core component of TPO) and NPO methods. “F” denotes forget quality, and “MU” denotes model utility.
Benchmark Used LLM Forget quality Model Utility
MUSE ICLM-7B KnowMem on $\mathcal{D}_{f}$ $\downarrow$ KnowMem on $\mathcal{D}_{r}\quad\uparrow$
VerbMem on $\mathcal{D}_{f}$ $\downarrow$
LLaMa-2 7B PrivLeak $(\rightarrow 0)$
TOFU LLaMa-2 7B Truth Ratio on $\mathcal{D}_{f}$ $\uparrow$ $\text{Mean}\Bigg(\begin{cases}\text{Probability, Rouge-L, Truth Ratio}\\ \mathcal{D}_{r},\mathcal{D}_{\text{real\_authors}},\mathcal{D}_{\text{word\_facts}}\end{cases}\!\!\!\!\!\!\!\Bigg)$ $\uparrow$
LLaMa-3.2 3B
Appendix C Unlearning performance combining the GDR with the proposed TIF framework.
Table 6 summarizes the forget quality and model utility of our TPO compared to several baseline approaches (i.e., GA, KTO, NPO, SimNPO), evaluated on the TOFU Forget 05 task. The experiments are conducted under two conditions: with and without incorporating Gradient Descent on Retain (GDR) loss, and with and without our GPT-based unwanted information identifier (GPT). Here, ”Vanilla” denotes methods evaluated without incorporating the GPT-based identifier.
@@
Appendix D Generative LM Approach vs Discriminative LM Approach.